OpenStudy (anonymous):

cosx+1/sin^3x=cscx/1-cosx

OpenStudy (loser66):

\(cos x + \dfrac{1}{sin^3 x}= \dfrac{csc x}{1-cos x}\) or \(\dfrac{cos x+1}{sin^3 x}=...\) ??which one?